OpenStudy (anonymous):

A circle is described by the equation (x−1/2 )^2 +(y−3/2 )^2 =4/9 . What are the cordinates for the center of the circle and the length of the radius? A) (-1/2 ,-3/2 ) , 2/3 units B) (-1/2 ,-3/2 ) , 4/9 units C) (1/2 ,3/2 ) , 4/9 units D) (1/2 ,3/2 ) , 2/3 units

OpenStudy (austinl):

I'm sorry, you need to clean up the presentation a little bit. As is, I am unable to tell just, what exactly, it is saying.

OpenStudy (anonymous):

BETTER?

OpenStudy (austinl):

Much better actually. Now, gimme a sec to think it over and find you a formula :)

OpenStudy (anonymous):

thank you :)

OpenStudy (austinl):

\((x-h)^2+y-k)^2=r^2\) \((h,k)\) is your center. \(r\) is your radius Now from what you can see here, what do you think would be the best answer? Sorry about the delay, I had to search like a mad man for my charger cable.

OpenStudy (anonymous):

b?

OpenStudy (austinl):

That is not the answer that I would go with. Try again :)

OpenStudy (anonymous):

Remember the signs of h and k change when you plug them into the equation @lindsay008

OpenStudy (anonymous):

d?

OpenStudy (austinl):

I would agree. \(\Large{\sqrt{\frac{4}{9}}=\frac{\sqrt{4}}{\sqrt{9}}=\frac{2}{3}}\) That is how the radius is derived as well, in case you just guessed :P

OpenStudy (anonymous):

so it should be d ? lol

OpenStudy (austinl):

Yes.

OpenStudy (austinl):

My question is this though, do you see how?